OpenStudy (anonymous):

Find the exact value of cos 11π/4

OpenStudy (anonymous):

Now pi/4 = 180/4 = 45 degrees. So we want to find cos of 45x11 = 495 . So are you aware that every 180 degrees the value of cos equals cos 0 and becomes 1?

OpenStudy (anonymous):

no :S but I am now

OpenStudy (anonymous):

ok sorry its 360 degrees; nt 180. :P Every 360 degrees 1 complete rotation is done; so cos 360 = cos 0 actually. So we're left with the other 495-360 = cos 135. If you want to know, then cos x = cos(180-x) => cos 135 = cos(180-135) => cos 135 = cos 45. However the angle of 135 lies in the II quadrant where cos is negative. So cos 45 = 1/ sqrt 2, however cos 135 will be -1/sqrt 2 also, cos 135 = cos of 3pi/4. = -1/sqrt 2

OpenStudy (anonymous):

got it! thanks :)

OpenStudy (anonymous):

alternatively, could i do this? 11π/4 = 12π/4 - π/4 = 3π - π/4 = 2π + π - π/4 2π = 0 (2π is one complete revolution) therfore, 11π/4 = π - π/4 in the 2nd quadrant, angles are in the form π-theta, so the angle is in the 2nd quadrant, and cos theta is neg in the 2nd quadrant therfore, cos 11π/4 = cos (π- π/4) = - cos π/4 = - 1/ sqrt 2

OpenStudy (anonymous):

the last step you did it in radians i did it in degrees. but yeah thats fine. :)

OpenStudy (anonymous):

ok, thanks